# Noncommutative Phantom BTZ Black Hole

### v1

@@ 4 Black hole as heat engine
Now, we analyze the black hole as a thermo-gravitational heat engine. A heat engine operates in a closed cycle within the $p-V$ plane, absorbing heat energy $Q_{H}$ from a high-temperature source, converting part of this energy into work $W$, and transferring the remaining heat $Q_{C}$ to a low-temperature reservoir. The efficiency $\Gamma$ of such an engine is defined as the ratio of the work performed to the heat absorbed, expressed as:
@@
To model the NC Phantom BTZ black hole as a heat engine, we consider a rectangular thermodynamic cycle in the $P-V$ plane, consisting of two isobaric and two isochoric processes, as illustrated in Figure 10.
Figure 10: Schematic representation of the heat engine cycle in the $P-V$ plane.
@@ 5 Conclusion
In this work, we have explored the thermodynamic and geometric properties of phantom BTZ black holes within the framework of noncommutative geometry. By incorporating Lorentzian distributions for mass and charge, we derived a modified metric that encapsulates the effects of non-commutativity on the spacetime structure. The analysis revealed significant alterations in horizon configurations, curvature singularities, and thermodynamic quantities due to the interplay between non-commutativity and phantom field dynamics.
Through a detailed and comparative thermodynamic investigation, we calculated key quantities such as the Hawking temperature, entropy, and heat capacity. The results indicate that noncommutative corrections profoundly influence the stability and phase transition behavior of phantom BTZ black holes. Notably, the noncommutative parameter introduces additional contributions to entropy and modifies the stability conditions, reflecting the impact of spacetime fuzziness on black hole thermodynamics. Furthermore, the investigation of the heat capacity revealed critical points associated with phase transitions, providing insights into the local and global stability of these systems.
The study also evaluated the black hole’s performance as a thermodynamic heat engine, highlighting the intricate role of noncommutative corrections in determining its efficiency. The findings demonstrate that the energy extraction potential of the black hole is sensitive to variations in the noncommutative parameter, which can enhance or suppress its efficiency. This analysis underscores the broader implications of noncommutative geometry for practical thermodynamic systems, offering a novel perspective on the interplay between quantum spacetime effects and gravitational thermodynamics.
Overall, this work emphasizes the significant role of noncommutative geometry in shaping the physical and thermodynamic properties of black holes, particularly when coupled with exotic fields. Future research could extend these findings by exploring higher-dimensional analogs of noncommutative phantom black holes or by incorporating additional fields to examine their combined effects. Moreover, numerical simulations of critical phenomena and the development of experimental analogs may provide valuable insights into the observable consequences of noncommutative geometry and exotic matter in gravitational systems.

### v2

@@ 4 Black hole as heat engine
Now, we analyze the black hole as a thermo-gravitational heat engine. A heat engine operates in a closed cycle within the pVp-V plane, absorbing heat energy $Q_{H}$ from a high-temperature source, converting part of this energy into work $W$, and transferring the remaining heat $Q_{C}$ to a low-temperature reservoir. The efficiency $\Gamma$ of such an engine is defined as the ratio of the work performed to the heat absorbed, expressed as:
@@
To model the NC Phantom BTZ black hole as a heat engine, we consider a rectangular thermodynamic cycle in the PVP-V plane, consisting of two isobaric and two isochoric processes, as illustrated in Figure 10.
Figure 10: Schematic representation of the heat engine cycle in the PVP-V plane.
@@ 5 Conclusion
In this study, we have examined the thermodynamic and geometric properties of phantom BTZ black holes in a noncommutative (NC) spacetime background. Using Lorentzian-type smeared energy-momentum tensors, we derived a novel modified BTZ metric that captures the combined effects of noncommutativity and exotic phantom fields. This approach offers a technically robust framework that regularizes singularities and incorporates quantum-spacetime features through a geometric lens.
Our investigation revealed several noteworthy results. First, we analytically computed key thermodynamic quantities—Hawking temperature, entropy, and heat capacity—under the influence of NC and phantom corrections. The noncommutative deformation introduces additional terms in the thermodynamic expressions, leading to modified entropy scaling and stability behavior. We identified critical points associated with second-order phase transitions and demonstrated that both local and global thermodynamic stability are sensitive to the NC parameter and phantom coupling.
Second, we formulated and analyzed the behavior of the black hole as a heat engine within the extended phase space formalism. We quantified the mechanical work output and efficiency for cyclic processes, showing that the energy extraction capability is significantly influenced by noncommutative and phantom effects. The introduction of NC geometry leads to nontrivial modulations in the PPVV diagram and alters the thermodynamic trajectories, thereby refining the standard heat engine framework.
The novelty of this work lies in the combined incorporation of Lorentzian noncommutative geometry and phantom matter within the BTZ black hole background, resulting in new insights into metric structure, phase behavior, and energy conversion. To our knowledge, this is one of the first comprehensive studies to analyze these two effects together in the context of $(2+1)$-dimensional black hole thermodynamics and heat engine efficiency.
Looking ahead, this framework can be generalized to include electric or magnetic charges, rotation, or higher-curvature corrections such as Gauss–Bonnet or Lovelock terms. Moreover, the thermodynamic geometry and holographic dual descriptions of such NC–phantom systems remain open avenues for exploration. Incorporating numerical analysis of critical exponents and exploring analog models may also help connect these theoretical insights with experimentally motivated contexts.